cash 25,000 shares of $18 par preferred stock at $61. a. Journalize the entries for October 31 and November 19. For a compound transaction, if an amount box does not require an entry, leave it blank. Oct. 31 Cash Common Stock Nov. 19 Cash Preferred Stock b. What is the total amount invested (total paid-in 929 AM
Answer #1
Dear Student
Thank you for using homeworklib
Please find below the answer
Statementshowing Computations
Oct 31 Debit Credit
Cash Dr = 160,000*12         1,920,000.00
To common stock= 160,000*8        1,280,000.00
To paid in capital in excess of par - Common Stock = 160,000*4            640,000.00
november 19
Cash Dr = 25000*61         1,525,000.00
To Preferred stock= 25000*18            450,000.00
To paid in capital in excess of par - preferred Stock = 25000*43        1,075,000.00
Nov 19
Total paid in capital = 1525,000 + 1920,000         3,445,000.00